The joint probability mass function of (X, Y) is given by p(x, y) = k(2x + 3y),

x = 0, 1, 2; y = 1, 2, 3. Find all the marginal and conditional probability

distributions. Also find the probability distribution of (X + Y)


1
Expert's answer
2021-06-29T03:57:41-0400
"\sum_x\sum_yp(x,y)=1""\begin{matrix}
 & & & y \\
 & p(x, y) & 1 & 2 & 3 \\
\hdashline 
 & 0 & 3k & 6k & 9k \\
 x & 1 & 5k & 8k & 11k \\
 & 2 & 7k & 10k & 13k \\
\end{matrix}"

"3k+6k+9k+5k+8k+11k+7k+10k+13k=1"

"k=\dfrac{1}{72}"

"x=0: p_X(0)=3k+6k+9k=18k=\dfrac{1}{4}"


"x=1: p_X(1)=5k+8k+11k=24k=\dfrac{1}{3}"

"x=2: p_X(2)=7k+10k+13k=30k=\dfrac{5}{12}"

"p_X(x)=\begin{cases}
 1/4 &\text{if } x=0 \\
 1/3 &\text{if } x=1 \\
 5/12 &\text{if } x=2 \\
 0 &\text{otherwise } \\
\end{cases}"

"y=1: p_Y(1)=3k+5k+7k=15k=\dfrac{5}{24}"


"y=2: p_Y(2)=6k+8k+10k=24k=\dfrac{1}{3}"

"y=3: p_Y(3)=9k+11k+13k=33k=\dfrac{11}{24}"


"p_Y(y)=\begin{cases}
 5/24 &\text{if } y=1 \\
 1/3 &\text{if } y=2 \\
 11/24 &\text{if } y=3 \\
 0 &\text{otherwise } \\
\end{cases}"


"p_{X|Y}(x|y)=\dfrac{p(x, y)}{p_X(x)}"

Conditional pmf's of "X"given "Y"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 x & p_{X|Y}(x|1) & p_{X|Y}(x|2) & p_{X|Y}(x|3) \\ \hline
 0 & 1/6 & 1/3 & 1/2 \\
 \hdashline
 1 & 5/24 & 1/3 & 11/24 \\
 \hdashline
 2 & 7/30 & 1/3 & 13/30 \\
 \hdashline
\end{array}"


"p_{Y|X}(yx)=\dfrac{p(x, y)}{p_Y(y)}"

Conditional pmf's of "Y"given "X"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 y & p_{Y|X}(y|0) & p_{Y|X}(y|1) & p_{Y|X}(y|2) \\ \hline
 1 & 1/5 & 1/3 & 7/15 \\
 \hdashline
 2 & 1/4 & 1/3 & 5/12 \\
 \hdashline
 3 & 3/11 & 1/3 & 13/33 \\
 \hdashline
\end{array}"


"x+y=1, (0,1):p(1)=3k=\dfrac{1}{24}"

"x+y=2, (0,2), (1, 1):p(2)=6k+5k=\dfrac{11}{72}"

"x+y=3, (0,3),(1,2),(2,1):p(3)=9k+8k+7k=\dfrac{1}{3}"

"x+y=4, (1,3),(2,2):p(4)=11k+10k=\dfrac{7}{24}"

"x+y=5, (2,3):p(15)=13k=\dfrac{13}{72}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x+y & 1 & 2 & 3 & 4 & 5 \\ \hline
 p(x+y) & 1/24 & 11/72 & 1/3 & 7/24 & 13/72 
\end{array}"